OpenStudy (anonymous):

There are 10 cars at the mechanic's garage. Each car needs 1 1/5 liters of oil. How much oil will the mechanic need for the cars? Express your answer in simplest form.

OpenStudy (anonymous):

@kittenlover731

OpenStudy (jojokiw3):

You have 10 cars. Each car needs 1 1/5 liters of oil. \[cars(liters) = total. \] So what you want to do is convert 1 1/5 into 6/5 \[cars(\frac{ 6 }{ 5 }) = total. \] Then plug in the number of cars. \[10(\frac{ 6 }{ 5 }) = ?\] Simply multiply THEN simplify and you have the answer. :)

OpenStudy (unanimoose):

1/5 = 0.2 1.2 * 10 = ?

OpenStudy (anonymous):

well, to simplify here is a calculator :) http://www.calculatorsoup.com/calculators/math/fractionssimplify.php

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

yep